# BIMCV-R: A Landmark Dataset for 3D CT Text-Image Retrieval

### v1

@@ Abstract
The burgeoning integration of 3D medical imaging into healthcare has led to a substantial increase in the workload of medical professionals. To assist clinicians in their diagnostic processes and alleviate their workload, the development of a robust system for retrieving similar case studies presents a viable solution. While the concept holds great promise, the field of 3D medical text-image retrieval is currently limited by the absence of robust evaluation benchmarks and curated datasets. To remedy this, our study presents a groundbreaking dataset, BIMCV-R [^1] This dataset will be released upon acceptance., which includes an extensive collection of 8,069 3D CT volumes, encompassing over 2 million slices, paired with their respective radiological reports. Expanding upon the foundational work of our dataset, we craft a retrieval strategy, MedFinder. This approach employs a dual-stream network architecture, harnessing the potential of large language models to advance the field of medical image retrieval beyond existing text-image retrieval solutions. It marks our preliminary step towards developing a system capable of facilitating text-to-image, image-to-text, and keyword-based retrieval tasks.
@@ 1 Introduction
The rapid evolution of medical imaging technologies, especially in the realm of 3D imaging, has brought about a transformative change in radiological diagnostics [31, 2, 13]. These technologies provide detailed, three-dimensional visualizations that are crucial for accurate lesion detection and measurement, which are pivotal in disease staging, treatment planning, and prognosis assessment [7, 23]. However, the sheer volume and complexity of this spatial data have substantially increased the workload of clinicians, prompting the need for Artificial Intelligence (AI) to assist in the analysis process. AI applications such as automatic segmentation [11, 4, 29, 16, 33, 14], reconstruction [25, 5, 17], and denoising [10, 30, 22] are now integral in providing clearer insights and improving diagnostic accuracy.
AI’s role in diagnostics is further expanded through image recognition tools, including detection [9, 15] and classification [8, 20], which are increasingly being used in conjunction with clinical textual reports to provide a more holistic understanding of medical conditions [18, 19, 24, 3]. Initiatives such as Quilt-1m [12], and pre-training projects like BiomedCLIP [32] and MedClip [27], as well as efforts in 3D image description by GTGM [6], T3D [17], and 3d-MIR [1], reflect the ongoing efforts to connect medical images with their textual counterparts. Despite these advancements, the focus on 2D image-text pairs and the challenges with model-generated descriptions highlight the need for a comprehensive benchmark that integrates 3D medical imaging with textual diagnostics, a significant void in the current landscape [28].
@@ 2 Dataset
This paper presents the BIMCV-R dataset, a substantial resource meticulously crafted for 3D medical multimodal retrieval. This dataset is an extension of the BIMCV dataset [26], encompassing pristine CT scan images, detailed radiological reports, and comprehensive DICOM metadata. We have secured permission from the BIMCV team and are committed to releasing the dataset to the public upon the publication of this paper.
@@
Table 1: Summary of Image and Report Statistics. Statistic Average Median Range Image Width 529 512 514 $\sim$ 710 Image Height 528 512 520 $\sim$ 672 Number of Slices 279 224 101 $\sim$ 670 Length of Report 104 97 7 $\sim$ 260
@@ 3 Methodology
3.0.3 Visual Feature Extraction.
@@
where $L_{sim}$ is the loss function, encouraging the model to increase the similarity score of positive pairs while decreasing that of negative pairs.

### v2

@@ Abstract
The burgeoning integration of 3D medical imaging into healthcare has led to a substantial increase in the workload of medical professionals. To assist clinicians in their diagnostic processes and alleviate their workload, the development of a robust system for retrieving similar case studies presents a viable solution. While the concept holds great promise, the field of 3D medical text-image retrieval is currently limited by the absence of robust evaluation benchmarks and curated datasets. To remedy this, our study presents a groundbreaking dataset, BIMCV-R, which includes an extensive collection of 8,069 3D CT volumes, encompassing over 2 million slices, paired with their respective radiological reports. Expanding upon the foundational work of our dataset, we craft a retrieval strategy, MedFinder. This approach employs a dual-stream network architecture, harnessing the potential of large language models to advance the field of medical image retrieval beyond existing text-image retrieval solutions. It marks our preliminary step towards developing a system capable of facilitating text-to-image, image-to-text, and keyword-based retrieval tasks. Our project is available at https://huggingface.co/datasets/cyd0806/BIMCV-R.
@@ 1 Introduction
The rapid evolution of medical imaging technologies, especially in the realm of 3D imaging, has brought about a transformative change in radiological diagnostics [1, 2, 3, 4]. These technologies provide detailed, three-dimensional visualizations that are crucial for accurate lesion detection and measurement, which are pivotal in disease staging, treatment planning, and prognosis assessment [5, 6]. However, the sheer volume and complexity of this spatial data have substantially increased the workload of clinicians, prompting the need for Artificial Intelligence (AI) to assist in the analysis process. AI applications such as automatic segmentation [7, 8], reconstruction [9, 10, 11, 12], and denoising [13] are now integral in providing clearer insights and improving diagnostic accuracy.
AI’s role in diagnostics is further expanded through image recognition tools, including detection [14] and classification [15], which are increasingly being used in conjunction with clinical textual reports to provide a more holistic understanding of medical conditions. Initiatives such as Quilt-1m [16], and pre-training projects like BiomedCLIP [17] and MedClip [18], as well as efforts in 3D image description by GTGM [19], T3D [20], and 3d-MIR [21], reflect the ongoing efforts to connect medical images with their textual counterparts. Despite these advancements, the focus on 2D image-text pairs and the challenges with model-generated descriptions highlight the need for a comprehensive benchmark that integrates 3D medical imaging with textual diagnostics, a significant void in the current landscape [22].
@@ 2 Dataset
This paper presents the BIMCV-R dataset, a substantial resource meticulously crafted for 3D medical multimodal retrieval. This dataset is an extension of the BIMCV dataset [23], encompassing pristine CT scan images, detailed radiological reports, and comprehensive DICOM metadata.
@@
Table 1: Summary of Image and Report Statistics.
Statistic Average Median Range
Image Width 529 512 514 $\sim$ 710
Image Height 528 512 520 $\sim$ 672
Number of Slices 279 224 101 $\sim$ 670
Length of Report 104 97 7 $\sim$ 260
@@ 3 Methodology
3.0.3 Visual Feature Extracting.
@@
where $L_{sim}$ is the loss function, encouraging the model to increase the similarity score of positive pairs while decreasing that of negative pairs [24].
@@ 5 Conclusion
Disclosure of Interests
The authors have no competing interests to declare that are relevant to the content of this article.